OpenStudy (anonymous):

Please help, I don't understand. If y varies directly as x, and y is 400 when x is r and y is r when x is 4, what is the numeric constant of variation in this relation?

OpenStudy (anonymous):

\[y=kx\] you need \(k\) you are given \[400=kr\] and \[r=4k\]

OpenStudy (anonymous):

that makes \(400=4k^2\) and so \(100=k^2\) making \(k=10\)

OpenStudy (anonymous):

Thanks so much. NOW i understand